1. Calculate the magnitude of the electric field at the center of a square 42.5cm on a side if one corner is occupied by a −38.9μC charge and the other three are occupied by −26.5μC charges. Choose the correct direction of the electric field at the center of the square.

2. Choose the correct direction of the electric field at the center of the square.

A. toward to −38.9μC charge

B. perpendicular to the diagonal passing through −38.9μC

C. charge outward of −38.9μC charge
Answer #1

Electric field at the center due to the charge 1
kqi
910 38.9 10- 0.302 E1
E1 = 3.89 N/C
Now the electric field due to the charge 2
kg2 E2 =下
F9+10 26.5+1 0.302
E2 = 2.65 N/C
Electric field due to the charge 3 and 4 will be same as E2 due to symmetry , therefore
E2 = E3 =E4
E2 and E4 will cancelled each other
Hence the net field at the center will be
E = E1 - E3 = 3.89 - 2.65 = 1.24 N/C
(2)
the direction will be toward - 38.9 muC
Hence the correct option is A. i.e toward to - 38.9 muC